# Query Rewriting via LLMs

### v1

@@ 4.1 Redundancy Removal
There are different types of redundancy that can occur in a SQL query – repeated computations, superfluous filter predicates, unnecessary joins, etc. Rules R1 through R4 in Table 2 are designed to tackle such redundancies. For instance, the redundant PK-FK join highlighted in the Introduction is addressed by rule R4 in Table 2 (the relevant schematic information is also provided in the prompt).
Table 2: Rules for Database-sensitive prompts
@@
R1 Use CTEs (Common Table Expressions) to avoid repeated computation.
@@
R6 Pre-filter fact tables in a CTE using dimension tables with low selectivities. Retain dimension table filters in main query. Do not create explicit join statements.
The template for such rule-based prompts is shown in Figure 5(a) and includes an example to demonstrate the rule application to the LLM – the specific examples used with our rules are enumerated in the Appendix A3. Note that this prompt template allows for only a single rule to be present in the prompt. This was a conscious design choice because LLMs are often overwhelmed by excessive information given in monolithic form. Therefore, we apply each rule using a separate prompt, finally returning the rewrite providing the best performance improvement.
@@
The performance improvement achieved on the micro-benchmark by an ensemble that adds the redundancy-removing prompts to the basic set (Section 3) is shown in Table 3. We observe that the PR increases to 8, and SpeedupGM grows to 4.8.
Table 3: Performance with Redundancy Removal Rules
Prompt # PR SpeedupGM
@@ 4.2 Selectivity-based Guidance
We now turn our attention to rules whose applicability to a query is conditional on the specific database environment, specifically its statistical aspects. For example, consider the alternative rewrites shown in Figure 6 using the EXIST and IN clauses (highlighted in red), respectively – here the appropriate choice is dictated by the selectivity of the inner subquery – EXISTS for high selectivity values and IN for low values. Based on such arguments, rules R5 and R6, which pre-filters fact tables for dimension tables with low selectivity values, are included in Table 2. Note that, a specific instruction to not create explicit joins had to be added in rule R6. This is because in the presence of CTEs, the LLM is prone to schematic confusion regarding which attribute belongs to which table, leading it to construct invalid joins.
@@
The performance improvements following addition of selectivity-guided prompts are shown in Table 4. We observe that PR are now obtained for all 10 micro-benchmark queries. Moreover, the resulting SpeedupGM increases to 11.8, quite close to the human upper bound of 13.9.
Table 4: Performance of Metadata-infused Prompts
Prompt # PR SpeedupGM
@@
A key point to note in closing is that rules R1 through R6 not only bring additional queries into the productive category, but also deliver greater improvement for those already deemed to be productive via the standard prompts of Section 3. Further, to minimize selection overheads, a classifier could be designed to choose the appropriate rule, and this option is examined in Section 7.3.3.
@@ 5 Token Probability Driven Rewrite
A key challenge when using LLMs is “hallucinations” – the generation of responses that range from being mildly incorrect to completely made up. This is often due to the potential output tokens having low confidence, which in turn results in “confusing” the LLM and generating suboptimal outputs. In order to have a robust approach for such cases, we take inspiration from the code generation literature [36]. Specifically, we propose a Monte Carlo Tree Search (MCTS) based decoding approach to search for a sequence of LLM-generated tokens that results in both a valid query rewrite as well as performance improvements.
@@
Given that the vocabulary sizes of LLMs are upward of hundreds of thousands of tokens, it may become very expensive (in both financial and computational costs) to construct the entire tree. It is therefore essential to significantly reduce the token search space while exploring the tree for valid rewrites. This is precisely the purpose of MCTS which applies an Upper Confidence Bound (UCB) heuristic [26] to identify the best paths in a tree without computing the entire tree.
5.1 MCTS Search Process
$n_{\text{next}}=\operatorname*{arg\,max}_{a\in A}UCB(n,a)$ (1)
$\beta(n)=log(\frac{visits[n]+c_{base}+1}{c_{base}})+c$ (3)
@@ 5.2 Input Prompt to MCTS
The root state in Algorithm 1 corresponds to the state just after the prompt is fed to the LLM. One way to use this algorithm is to execute it for all the various prompts discussed in the previous sections, and choose the rewrite that provides the best performance. This, however, is expensive both from the aspect of query rewrite time, as well as the number of LLM tokens used. To minimize these costs, the LITHE workflow first selects, given a query, the prompt yielding the most effective rewrite from among the techniques of Sections 3 and 4. It then employs this prompt to initiate the MCTS-based rewrite. In case no prompt provides a lower-cost rewrite, baseline Prompt 1 of Section 3 is used as the fallback option.
Later, in Section 7.3.3, we show how the overheads incurred can be further reduced via a classifier construction.
@@ 5.3 Performance
The performance results with MCTS-based rewrites are shown in Table 5 for the micro-benchmark. We observe gratifyingly that the SpeedupGM now attains the human speedup of 13.9.
Table 5: Performance of MCTS-based rewrites
# PR SpeedupGM
LITHE 10 13.9
@@ 6.1 LITHE Parameter Settings
The “temperature” parameter of GPT-4o, which ranges over [0,1], controls the randomness of the model’s response, with a higher temperature resulting in the outputs being more varied and less predictable. While a higher temperature can be useful for creative writing where one would seek diverse and exploratory answers, in our case we want a focused and deterministic answer as far as possible. Hence we set the temperature to 0 which causes the model to do greedy sampling to select the next token.
@@
Finally, we run the prompt pipeline for a maximum of 5 iterations to try and fix any syntax errors created during the rewrite process.
6.2 Query Efficiency Characterization
The LITHE workflow requires computing the performance of several candidate query rewrites. In this scenario, explicitly executing each rewrite on the database to evaluate its runtime benefit may become impractically slow. This is especially true for queries where even a single execution is time-consuming. We therefore leverage the query cost, as estimated by the optimizer, to evaluate the performance of query rewrites. It is common knowledge that these estimates are often inaccurate [16] – however, they are expected to be generally reflective of the actual run times, especially on modern industrial-strength database systems. Moreover, our comparison of costs is only in a relative sense – between the original query and its rewrite, so errors in the absolute values are tolerable as long as similar errors are made in all estimates. These costs can therefore serve as a computationally cheap and meaningful measure to distinguish performance among rewrites.
@@ 6.3 Query Equivalence Testing
1. Logic-based Equivalence. Although verifying the equivalence of a general pair of SQL queries is a long-standing hard problem [6], a variety of logic-based tools are available for proving equivalence over restricted classes of queries (e.g. Cosette[10], SQL-Solver [13], QED [30]). In our pipeline, we used the recently proposed QED [30] since it was found to cover a larger set of queries compared to the alternatives. The advantage of such a logic-based approach is that it is definitive in outcome and relatively inexpensive.
2. Result Equivalence via Sampling. If the original query is not within the QED scope, we alternatively use a sampling-based approach to test equivalence. The idea here is to execute the queries on a small sample of the database and verify equivalence based on the sample results. However, while this test is a necessary condition for query equivalence, it is not sufficient. That is, false positives may be present because the sampled database may not cover all the predicates featured in the query. To minimize this likelihood, we use a combination of (1) correlated sampling [34] for maintaining join integrity in the sample, (2) addition of synthetic tuples in the sample to distinguish outer and inner joins, and (3) adjusting constants in the filter predicates to produce populated results – the complete details are in the Appendix A2.
3. Result Equivalence on the Entire Database. Instead of resorting to samples, result equivalence could also be evaluated, in principle, on the entire database itself. However, as mentioned above, this could turn out to be prohibitively expensive, especially if the queries themselves are time consuming (e.g., due to the scale of the underlying database) and/or if the candidate rewrites happen to be regressions. Therefore, this option is disabled by default in our workflow, but the user can optionally invoke it as a final verification step after the sampled equivalence testing.
@@ 7 Experimental Evaluation
In this section, we report on LITHE’s performance profile. We first describe the experimental setup, including comparative baselines, query suites and evaluation platforms. Then we present the speedup results for both aggregate benchmark and individual queries, followed by characterization of the rewrite overheads in computational and financial terms. We finally discuss the impact of alternative platforms wrt database schema, database engine and LLMs.
@@
Given an input query, each of the SOTA approaches is independently invoked to perform a rewrite, and the rewrite with the best performance is used as the baseline for comparison. Note that these approaches may generate rewrites that regress the performance [19]. We discard such backsliding rewrites using, similar to LITHE, the Query-Costing module of the database engine.
Query Set. Our evaluation uses the FPR subsets from the following diverse set of benchmarks, amounting to 100+ queries:
  1. 1.
    TPC-DS [11]: A synthetic OLAP benchmark modeling retail supplier environment with complex analytical queries.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
@@
Testbed. The majority of our experiments are carried out on the following data processing platform: Sandbox server with Intel(R) Xeon(R) CPU E5-1660 v4 @ 3.20GHz x 16, 32 GB RAM, and 12TB HDD, running Ubuntu 22.04 LTS; PostgreSQL v16 database engine; and GPT-4o LLM for both LITHE and SOTA. Variations on this platform are considered in Section 7.4.
Metrics. For each benchmark and each rewrite technique, we identified the number of queries for which a PR (productive rewrite) could be constructed. Subsequently, we computed the GM of the speedups obtained by each technique over the set of FPR (feasible productive rewrites). Finally, recall that if a rewrite achieves the maximum known speedup for the query, it is referred to as an MPR (Maximum Productive Rewrite).
@@
From the investment perspective, we measured the average rewrite time per query, and for the LLM-based techniques, the number of tokens used in the rewrite process.
7.1 Rewrite Quality
Table 6: Comparing LITHE with SOTA on FPR queries
Benchmark (FPR) # PR SpeedupGM
LITHE SOTA LITHE SOTA
TPC-DS (65) 65 16 30.6 2.1
DSB (20) 20 3 22.6 1.3
ARCHER (25) 25 19 2.2 1.8
JOB (5) 5 2 1.8 1.3
StackOverflow (2) 2 1 9.0 7.5
7.1.1 Benchmarks
The speedup delivered by LITHE and SOTA on the FPR queries of the various benchmarks are shown in Table 6. The good news is that for TPC-DS and its DSB variant, LITHE produces PR for all the FPR queries – 65 in TPC-DS and 20 in DSB. In fact, most of these rewrites are highly productive as the SpeedupGM exceeds 30 for TPC-DS and 20 for DSB! Further, the PR coverage is far superior in comparison to SOTA– by a factor of 4 (TPC-DS) and 6 (DSB). Finally, the SpeedupGM of LITHE is higher by more than an order-of-magnitude wrt SOTA.
Figure 7: Rewritten TPC-DS Q88 produced by LITHE
7.1.2 Query Readability
We now turn our attention to assessing how the readability of the queries is impacted by rewrites. That is, we wish to test whether the common wisdom that “efficient queries are more readable” holds for our rewrites. For this purpose, the number of joins in the query is used as a rudimentary metric to quantify query readability. The intuition is that a query with more joins is harder for the user to parse and understand.
Table 7 presents the average number of joins in the original and rewritten queries for both SOTA and LITHE. The results clearly demonstrate that LITHE significantly reduces the number of joins, especially for TPC-DS, DSB and ARCHER benchmark, thereby enhancing query readability. Notably, for the DSB benchmark, SOTA actually worsened the readability, while LITHE maintained superior readability compared to the original query. However, for the JOB and StackOverflow benchmarks, where optimization opportunities were limited, the readability remained unchanged. Importantly, LITHE never performs worse than the original or SOTA-rewritten queries in terms of readability, always maintaining or improving the metric.
Figure 7 shows the rewritten query for Q88 in which the number of joins reduced from 24 to 3. Overall, LITHE proves to be a reliable and effective tool for query rewriting, delivering both better performance and improved readability.
Table 7: LITHE vs SOTA readability comparison
Benchmark (FPR) Average # JOINS
LITHE SOTA Original
TPC-DS (65) 4.1 4.4 4.4
DSB (20) 4.8 5.3 4.95
ARCHER (25) 0.8 1 1.44
JOB (5) 7 7 7
StackOverflow (2) 8 8 8
Turning our attention to the other benchmarks (ARCHER, JOB, StackOverflow), the number of FPR queries is smaller due to the predominance of flat SPJ formulations in these benchmarks, which limits the scope for productive rewriting. Nevertheless, LITHE continues to achieve PR in all of this limited set, whereas SOTA misses quite a few opportunities. Further, the SpeedupGM of LITHE is visibly better than SOTA.
Figure 8: Speedups achieved by LITHE and SOTA

### v2

@@ 4.1 Redundancy Removal
There are different types of redundancy that can occur in a SQL query – repeated computations, superfluous filter predicates, unnecessary joins, etc. Rules R1 through R4 in Table 2 are designed to tackle such redundancies. The relevant schematic information (e.g. table names, column names, constraints) required by these rules is also provided in the prompt.
Table 2: Rules for Database-sensitive prompts.
@@
R1 Use CTEs (Common Table Expressions) to avoid repeated computation of a given expression.
@@
R6 Pre-filter tables involved in self-joins and with low selectivities on their filter and/or join predicates. Remove any redundant filters from the main query. Do not create explicit join statements.
@@
The template for such rule-based prompts is shown in Figure 5(a) and includes an example to demonstrate the rule application to the LLM – the specific examples used with our rules are available in the Section E. Note that this prompt template allows for only a single rule to be present in the prompt. This was a conscious design choice because LLMs are often overwhelmed by excessive information given in monolithic form. Therefore, we apply each rule using a separate prompt, finally returning the rewrite providing the best performance improvement.
@@
The performance improvement achieved on the micro-benchmark by an ensemble that adds the redundancy-removing prompts to the basic set (Section 3) is shown in Table 3. We observe that the CPR increases to 7, and CSGM grows to 6.85.
Table 3: Performance with Redundancy Removal Rules.
Prompt # CPR CSGM
@@ 4.2 Selectivity-based Guidance
We now turn our attention to rules whose applicability to a query is conditional on the specific database environment, specifically its statistical aspects. For example, consider the alternative rewrites shown in Figure 6 using the EXIST and IN clauses (highlighted in red), respectively – here the appropriate choice is dictated by the selectivity of the inner subquery – EXISTS for high selectivity values and IN for low values. Rule R5 basically encodes this argument as a rule in Table 2. Similarly rule R6, which pre-filters tables that are involved in self-joins and have low selectivity filter and/or join predicates, is also used to guide a rewrite. Note that a specific instruction to not create explicit joins had to be added in rule R6. This is because in the presence of CTEs, the LLM is prone to schematic confusion regarding which attribute belongs to which table, leading it to construct invalid joins.
@@
The performance improvements following addition of selectivity-guided prompts are shown in Table 4. We observe that CPRs are now obtained for 9 of the ten micro-benchmark queries. Moreover, the resulting CSGM increases to 10.57, quite close to the human target of 11.84.
Table 4: Performance of Metadata-infused Prompts.
Prompt # CPR CSGM
@@
We note in closing that rules R1 through R6 not only add queries to the productive category, but also deliver greater improvement for those already deemed to be productive via the standard prompts of Section 3. Further, to minimize selection overheads, a classifier could be designed to choose the appropriate rule – this option is examined in the Section B.
@@ 5 Token Probability Driven Rewrite
A key challenge when using LLMs is “hallucinations” – the generation of responses that range from being mildly incorrect to completely made up. This is often due to the output tokens having low confidence, which in turn results in “confusing” the LLM and generating suboptimal outputs. In order to have a robust approach for such cases, we take inspiration from the code generation literature [43]. Specifically, we propose a Monte Carlo Tree Search (MCTS) based decoding approach to search for a sequence of LLM-generated tokens that results in both a valid query rewrite as well as performance improvements.
@@
Given the hundreds of thousands of tokens in LLM vocabularies, it may be very expensive (both financially and computationally) to construct the entire tree. It is therefore essential to significantly reduce the token search space while exploring the tree for valid rewrites. This is precisely the purpose of MCTS which applies an Upper Confidence Bound (UCB) heuristic [31] to identify the best paths in a tree without computing the entire tree.
@@
5.1 MCTS Search Process
$n_{\text{next}}=\operatorname*{arg\,max}_{a\in A}UCB(n,a)$ (1)
$\beta(n)=log(\frac{visits[n]+c_{base}+1}{c_{base}})+c$ (3)
@@ 5.2 Input Prompt to MCTS
The root state in Algorithm 1 corresponds to the state just after the prompt is fed to the LLM. One way to use this algorithm is to execute it for all the various prompts discussed in the previous sections, and choose the rewrite that provides the best performance. This, however, is expensive both from the aspect of query rewrite time, as well as the number of LLM tokens used. To minimize these costs, given a user query, the LITHE workflow first selects the prompt yielding the most effective rewrite from among the techniques of Sections 3 and 4. It then employs this prompt to initiate the MCTS-based rewrite. In case no prompt provides a lower-cost rewrite, baseline Prompt 1 of Section 3 is used as the fallback option.
@@ 5.3 Performance
We find that LITHE’s performance with MCTS-based rewrites fully matches the human target, providing 10 CPRs and the maximum CSGM of 11.84. While these improvements, as compared to just prompting, might seem marginal at this stage, we wish to make two observations: (1) MCTS helps extract the maximum improvement wherever possible; and (2) As shown later in Section 7.5.2, improvements due to MCTS become significant for smaller LLMs such as LLaMA.
@@ 6.1 LITHE Parameter Settings
The “temperature” parameter of GPT-4o, which ranges over [0,1], controls the randomness of the model’s response. While a higher temperature can be useful for creative writing where one would seek diverse and exploratory answers, in our case we want a focused and deterministic answer as far as possible. Hence we set the temperature to 0 which forces the model to greedily sample the next token.
@@
Finally, we try a maximum of 5 times to fix, via prompt corrections, any rewrite that exhibits syntax errors (Section 2).
@@ 6.2 Query Equivalence Testing
1. Logic-based Equivalence. Although verifying the equivalence of a general pair of SQL queries is NP-complete [6], a variety of logic-based tools (e.g. Cosette[11], SQL-Solver [15], VeriEQL [17], QED [36]) are available for proving equivalence over restricted classes of queries, as mentioned in the Introduction. In LITHE, we use the recently proposed QED [36] since it was found to cover a larger set of queries compared to the alternatives. The advantage of such a logic-based approach is that it is definitive in outcome and relatively inexpensive.
2. Result Equivalence via Sampling. If the original query is not within the QED scope, we alternatively use a sampling-based approach to test equivalence. The idea here is to execute the queries on several small samples of the database and verify equivalence based on the sample results. However, while this test is a necessary condition for query equivalence, it is not sufficient. That is, false positives may be present because the sampled database may not cover all the predicates featured in the query. To minimize this likelihood, we use a combination of (1) correlated sampling [42] for maintaining join integrity in the sample, (2) adding synthetic tuples in the sample to distinguish outer and inner joins, and (3) adjusting constants in the filter predicates to produce populated results – the complete details are in the Section C.
3. Result Equivalence on the Entire Database. Result equivalence could also be evaluated, in principle, on the entire database itself. Of course, this could turn out to be prohibitively expensive, especially if the queries themselves are time-consuming (e.g. due to the scale of the underlying database) and/or if the candidate rewrites happen to be regressions. Therefore, we leave this check as an optional choice for the DBA.
@@ 7 Experimental Evaluation
In this section, we report on LITHE’s performance profile. We first describe the experimental setup, including comparative baselines, query suites and evaluation platforms. Then we present the speedup results for both aggregate benchmark and individual queries, followed by characterization of the rewrite overheads in computational and financial terms. We finally discuss the impact of alternative platforms wrt database engine, database schema and LLMs.
@@
Given an input query, each of the SOTA approaches is independently invoked to perform a rewrite, and the rewrite with the best performance is used as the baseline for comparison. Note that these approaches may occasionally generate rewrites that are expected by the optimizer’s costing module to regress the performance. For safety, we immediately discard such rewrites, similar to LITHE.
Query Set. Our evaluation in this paper primarily focuses on complex analytical queries from the standard TPC-DS decision-support benchmark [12], which models a retail supplier environment. The benchmark is used at its default size of 100 GB. As mentioned earlier, we focus on “slow queries” that take over 10 seconds to execute in their original form, operating in a cold-cache environment.
LITHE has also been evaluated on other benchmarks, including DSB [13], ARCHER [44], JOB [19] and StackOverflow [27]. Due to space limitations, we defer their details to Section A.
@@
Testbed. The majority of our experiments were carried out on the following data processing platform: Sandbox server with Intel(R) Xeon(R) CPU E5-1660 v4 @ 3.20GHz x 16, 32 GB RAM, and 12TB HDD, running Ubuntu 22.04 LTS; PostgreSQL v16 database engine; and GPT-4o LLM for both LITHE and SOTA. Variations on this platform are considered in Sections 7.4 and  7.5.
Metrics. For each rewrite technique, we identified the number of queries for which a CPR (cost productive rewrite with ¿ 1.5 speedup) could be constructed. Subsequently, we computed the CSGM (Cost Speedup Geometric Mean) and TSGM (Time Speedup Geometric Mean) performance obtained by each technique over the set of all CPRs (i.e. CPRs arising from either LITHE or SOTA).
@@
From the investment perspective, we measured the average rewrite time per query, and additionally for the LLM-based techniques, the number of tokens used in the rewrite process.
7.1 Rewrite Quality (Cost and Time)
7.1.1 Estimated Cost
LITHE produces a rewrite with a positive speedup ($>$ 1x) for 46 of the 88 TPC-DS queries deemed to be slow by our threshold. Of these 46, there were 26 CPRs resulting in a highly productive CSGM of 11.5. On the other hand, SOTA delivers only 13 CPRs (out of 42 positive rewrites) with a CSGM of 6.1. All but one of the SOTA CPRs also feature in the LITHE CPRs, making the total number of CPRs considered to be 27.
We were able to prove the rewriting correctness of 9 CPRs through the logic-based QED tool. The remaining passed our result equivalence tests on both the down-sampled and full databases. Furthermore, as a final check, we manually verified the correctness of the rewritten queries.
A drill-down into the cost speedup performance at the granularity of individual queries is shown in Figure 8, which compares LITHE (orange bars) and SOTA (blue bars) on each of the 27 CPR queries – note that the cost speedups on the $x$-axis are tabulated on a $\mathbf{log_{10}}$ scale, and the queries are sequenced in decreasing order of LITHE speedup. The vertical dotted line at $1$ represents the normalized baseline cost of the original query with the native optimizer, while the vertical line at $1.5$ is the CPR threshold.
We first observe, gratifyingly, that rewrites are indeed capable of promising dramatic cost speedups – take, for instance, Q41, which improves by a whopping five orders-of-magnitude for both SOTA and LITHE. This improvement in query performance is due to replacing the “WHERE (SELECT COUNT(*) from ...) > 0” clause with “WHERE EXIST (SELECT 1 from ...)” – the latter is a more efficient check for result existence in an inner subquery since it removes the computationally expensive aggregation function.
Figure 7: Plan Cost Speedups
Figure 8: Execution Time Speedups
Second, in most queries, the cost speedup of LITHE either exceeds or matches SOTA. In fact, there are several queries (e.g. Q45, Q25, Q4) where LITHE produces a highly beneficial CPR but SOTA simply returns the original query. Conversely, the opposite is true for one query (Q57) where SOTA projects a large speedup but LITHE settles for the original query. And in Q88 and Q95, SOTA is marginally better than LITHE.
At this stage, one could imagine that more rules could be added to LITHE to bring it on par with SOTA for queries like Q57. However, as discussed later in Section 8.3, such customized rules, while attractive at compile-time, run a significant risk of causing regressions at run-time. We have therefore selected only broad-brush rules for inclusion in LITHE. This conservative approach is justified below in the timing section, where we find that LITHE actually performed better than SOTA on these queries (Q57, Q88, Q95) wrt runtimes.
7.1.2 Execution Time
Thus far, we had considered optimizer-estimated execution costs. We now move to wall-clock runtimes for query executions – Figure 8 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA. We observe, again gratifyingly, that there are indeed several queries where substantial time benefits are achieved by the rewrites, even exceeding order-of-magnitude benefits in some cases – for instance, LITHE improves Q45 by a huge factor of 10000! Second, in almost all cases, LITHE outperforms or matches SOTA, including as mentioned above, the queries where SOTA’s optimizer costs were better. There remain only a few instances (Q9, Q47) where SOTA is marginally better.
From a modeling perspective, we see that the well-documented gap between optimizer predictions and actual run-times is prevalent in the rewrite space as well. On the one hand, there is Q45 where the projected speedup of $10^{2}$ increases to a huge $10^{4}$ at runtime. On the other hand, the $10^{5}$ speedup for Q41 decreases to $10^{3}$. But for SOTA, the reductions can be severe – a striking case in point is Q57, where SOTA actually causes regression despite a speedup projection of close to 100. However, the good news is that with LITHE, although the runtime speedups did not always match the projections, we did not encounter any regressions among the CPR rewrites. Overall, LITHE produces more robust rewrites resulting in a high TSGM of 18.4, whereas SOTA provides a TSGM of 6.
7.1.3 Reasoning
As a confirmatory exercise, we compared the LLM-generated explanation output by LITHE with our own manual analysis of the query plans generated for the original and rewritten queries. For example, consider Q45 – the provided explanation correctly lists the main reasons for the huge $10^{4}$ speedup as “the use of CTEs”, “Reduction in data volume early on” through pre-filtering based on join predicates, etc. Overall, we found that the explanations provided by LITHE matched our manual analysis of the plans, indicating that model-based reasoning is well aligned with human-backed reasoning in these scenarios.
7.2 Ablation Analysis of Rewrite Quality
7.2.1 Components of LITHE
A natural question at this stage is the role of the various techniques in LITHE towards achieving its large performance benefits. This analysis is captured in Table 5 which lists the CPRs for each technique when invoked in isolation as well as the cumulative number of CPRs when the different techniques are combined in the order of listing.
Table 5: CPRs contributed by LITHE components.
# CPR
Isolated Cumulative
Basic Prompts (Section 3) 15 15
Rules R1 — R4 (Section 4.1) 10 18
Rules R5, R6 (Section 4.2) 11 25
MCTS (Section 5) 26 26
We observe that the Basic Prompt ensemble and Redundancy Rules (R1 – R4) contribute to around two-thirds of the CPRs (18/26). When the statistics-infused Rules (R5, R6) are added, this number jumps to 25. Finally, MCTS produces an additional CPR and also improves the cost speedup of a pre-existing CPR.
7.2.2 Database-Sensitive Prompts.
Since MCTS explores the search space at a fine granularity, one could ask whether just the Basic Prompts in conjunction with MCTS would suffice to provide good performance. The motivation is that it would relieve us from using the database-sensitive rules that incur significant computational and financial overheads. When this experiment was conducted, the CSGM dropped precipitously to a paltry 5, a far cry from the 11.5 obtained with the database-sensitive rules. These results highlight the need to reflect database awareness for effective query rewriting, and not rely solely on prior LLM knowledge.
7.3 Rewrite Overheads (Time/Money)
Having established the performance benefits of rewrites, we now turn our attention to their time and financial overheads.
7.3.1 Transformation Time
The average processing time per CPR query is shown in Table 6, and we see that they do take a few minutes. However, this investment may be acceptable in deployment given that the execution benefits typically far outweigh the compilation overheads. For instance, with Q11, the original query took nearly two hours to complete, whereas the LITHE rewrite executed in just 3 minutes. Further, many applications tend to use a set of canned queries which are run thousands of times. Thus, even a large one-time investment can be easily recovered over repeat executions of such queries.
Table 6: Rewrite Overheads of LITHE and SOTA.
Avg. Time (min) Avg. Tokens Avg. Cost (USD)
LITHE 5 18427 0.045
SOTA 1.7 20076 0.050
Notwithstanding the above, we also observe that LITHE is considerably slower than SOTA in producing rewrites. A drill-down shows that the lion’s share of the time is taken by the initial prompt ensemble and the final MCTS module. We explore techniques to improve their efficiency in the Section B.
7.3.2 Monetary Outlay
The average number of LLM tokens required by LITHE and SOTA, and their associated financial costs [^1] At the time of writing, GPT-4o costs USD 2.5 per million tokens., are also shown in Table 6. The good news is that the inference charges per query are just a few cents, making rewriting practical from a deployment perspective.
7.4 Commercial DBMS
A legitimate question could be whether the rewrites made amends for the PostgreSQL optimizer but may fail to be useful in highly-engineered database engines. To evaluate this issue, we performed TPC-DS rewrites on a pair of popular commercial DBMS, which we refer to as OptA and OptB.
Table 7: Performance on Commercial Database Engines.
# CPR CSGM TSGM
OptA OptB OptA OptB OptA OptB
LITHE 7 8 3.6 4.8 1.5 2.4
SOTA 3 6 2.1 3.5 1.3 1.7
The performance of LITHE and SOTA on these two systems is shown in Table 7, with LITHE continuing to do better than SOTA. Despite the apparent lack of optimization headroom, LITHE still produces 7 and 8 CPRs resulting in a healthy CSGM of 3.6 and 4.8, respectively. Further, the TSGM provided by these rewrites are a useful 1.5 and 2.4, respectively. However, unlike the no-regression outcome seen with PostgreSQL, LITHE generates 3 CPRs on OptA, and 2 on OptB, which turn out to be regressions. So, although these commercial optimizers may have greater sophistication, they also suffer significant brittleness with regard to estimation accuracy.
Interestingly, one regression on OptB was caused by the LLM bringing in, via a basic prompt, an unrestricted version of Rule 6, which is prone to generating poorly-estimated plans (see Section 8.3). Despite such occasional errors, LITHE provides good overall improvement in runtimes as indicated by the TSGM values.
We also performed a preliminary study of LITHE on an internal real-world benchmark, whose data and query characteristics are more realistic than TPC-DS. Even on this hard benchmark, LITHE produced 8 CPRs (out of 45 queries tested) with a CSGM of 2 and a TSGM of 3.3.
The above results suggest LITHE has a useful role to play in industrial environments. From a different perspective, a company building a new database engine could use LITHE to non-invasively overcome the limitations of early versions of its optimizer.
7.5 Dependence on LLM (Training/Model)
7.5.1 Masked Database
An interesting question to ask now is whether the performance benefits seen thus far could be an artifact of GPT-4o having already been trained well on the TPC-DS benchmark, which is prominent in the public domain. To investigate this issue, we created a masked version of the TPC-DS database schema, whereby the table and column names convey no semantic information about their contents. We then constructed rewrites for the CPR queries (after syntactic changes to reflect the new masked schema) on this version. The results are shown in Table 8, and we observe that the performance profiles only marginally decrease for both LITHE and SOTA, testifying to their robust applicability.
Table 8: Rewrite Performance on Masked Database.
Approach # CPR CSGM
TPC-DS Masked TPC-DS Masked
SOTA 15 12 6.1 5.3
LITHE 26 24 11.5 10.5
7.5.2 LITHE on LLaMA
In our concluding experiment, we evaluated LITHE’s performance on the LLaMA 3.1 70 billion parameter instruct model, substantially smaller compared to GPT-4o, which may have several hundred billion parameters [5]. To attain practical inference times, the model was loaded using a low 4-bit quantization. Further, to ensure reproducibility and deterministic answers, the do_sample parameter was set to False, which forces the LLM to perform greedy decoding. To make up for the huge reduction in model parameters as compared to GPT-4o, we include up to two example demonstrations for each rule-based prompt.
For this environment, Table 9 shows LITHE’s performance on the TPC-DS benchmark with and without MCTS. Although certainly lower than the corresponding numbers with GPT-4o (Section 7.1), it is encouraging to see that, in absolute terms, significant performance benefits can be obtained for most queries, especially with MCTS support. So, the message is that smaller models can also be fruitfully used in real-world environments.
Table 9: LITHE Rewrite Performance with LLaMA.
# CPR CSGM TSGM
LLaMA without MCTS 17 5.6 8.2
LLaMA with MCTS 21 8.3 14.7
8 The Road Ahead
Based on our study, we now present a few observations with implications for the future design and deployment of rewriting tools.
8.1 Rewrite Space Coverage by LLMs
Given the decades-long research on database query optimization, we expected the potential for performance improvement via rewriting to be limited. What came as a surprise was the substantial scope for improvement still available, as showcased by the large CSGM and TSGM values, even on commercial platforms. These results suggest that LLMs explore optimization spaces that are well outside the purview of contemporary database engines. Further, this enhanced space could be augmented, in a two-stage process, with the recent proposals for LLM-based “plan hints” that steer the optimizer in fruitful directions within a plan space [7].
8.2 Rewrite Migration to Optimizer
The above demonstrated the potent exploratory power of LLMs. But from an overheads perspective, such rewrites should ideally be within the optimizer’s native search space rather than recommended from outside. Therefore, it would be a useful exercise to try and distill fresh optimization rules from these instances, leveraging the extensibility features of contemporary optimizers [14] to facilitate their incorporation in existing systems.